Z is the standard normal variable find P(-3.50<z<1.000)
Answer #1

We know that

Z is the standard normal variable

Z ~ Norm ( 0,1)

By using Z table (probability table )

P ( -3.50 < z < 1 ) = P (z<1) - P (z < -3.50)

= P (z<1) - [ 1 - P (z<3.50) ]

= 0.8413 - ( 1 - 0.9998 )

= 0.8411